OpenStudy (anonymous):

medal and fan!!!!! Two lines, A and B, are represented by the equations given below: Line A: 2x + 2y = 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (8, 4). It is (4, 8). There is no solution. There are infinitely many solutions.

OpenStudy (anonymous):

A student is trying to solve the set of two equations given below: Equation A: x + z = 6 Equation B: 2x + 3z = 1 Which of the following is a possible step used in eliminating the z-term? Multiply equation B by 3. Multiply equation A by 2. Multiply equation B by 2. Multiply equation A by −3.

OpenStudy (perl):

hint -3z + 3z = 0

OpenStudy (anonymous):

i don't understand?

OpenStudy (perl):

if you multiply equation A by -3, then add to equation B , what happens

OpenStudy (anonymous):

im trying to do it and i cant get the same equation as you

OpenStudy (anonymous):

if you multiply equation A by -3, then add to equation B. how do you get -3z + 3z = 0

OpenStudy (hockeychick23):

Equation A: x + z = 6 Equation B: 2x + 3z = 1 Which of the following is a possible step used in eliminating the z-term? Multiply equation B by 3. Multiply equation A by 2. Multiply equation B by 2. Multiply equation A by −3. So you don't care about what can get rid of the x-term, therefore you can eliminate answers B and C. Do you get that part so far?

OpenStudy (hockeychick23):

no it isn't

OpenStudy (anonymous):

yeah i got it

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

right?

OpenStudy (hockeychick23):

yepp thats right :) good job

OpenStudy (anonymous):

wait -3(3) equals -9

OpenStudy (anonymous):

and can you help with the second one please

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

sorry the second one is D

OpenStudy (anonymous):

thats okay thnx, can youy help with the first one

OpenStudy (anonymous):

I think is D because these two lines actually are the same, technically just the equation x+y=4 then x and y could be everything as long as their sum is 4. But I am not sure...Sorry